OpenStudy (anonymous):

please help will give medal Solve by completing the square: x2 + 2x - 24 = 0 A. -6, 4 B. 6, -4 C. 6, 4 D. 2, -3

jhonyy9 (jhonyy9):

what you ned to complete it ? you can this factorizing using these choises can you make it ?

jhonyy9 (jhonyy9):

B will be right because you can factorizing this using 6 and -4 (x+6)(x-4) = x^2 +2x -24

OpenStudy (anonymous):

thanks

jhonyy9 (jhonyy9):

yw

OpenStudy (whpalmer4):

To complete the square, put it in this form: \[x^2+2x=24\] (all x terms on 1 side, all numbers on other) Next, take half the value of the coefficient of the x term, square it, and add that quantity to both sides. \[x^2+2x+(2/2)^2 = 24 + (2/2)^2\] Now you can write the left side as a perfect square: \[(x+1)^2 = 25\] Solve that and you get your two answers in the way the you are supposed to...

OpenStudy (whpalmer4):

(remember that \(u^2 = 25\) has two solutions, \(u = \pm 5\))